Hello. Can you please help me solve these problems?
2x-y=-4
-4x+2y=6
and
2t+4v=6
-t-2v=-3
Thank you!

Answer by checkley77(12844)

2x-y=-4 MULTIPLY BY 2 & ADD.
-4x+2y=6
4x-2y=-8
------------
0x+0x=-2 NO unique solution for x or y.
2t+4v=6
-t-2v=-3 multiply by 2 & add
-2t-4v=-6
2t+4v=6
--------------
0t+0t=0 NO unique solution for x or y.